Buzz's Educational Software Outlet sells two or more of the video games as a single package. Managers are keenly interested in individual product-profitability figures. Information pertaining to three bundled products and the stand-alone prices is as follows:

Stand-Alone Selling Price

Cost

Package

Packaged

Price

Reading Fun

$20

$3.75

1. Reading Fun & Math Fun

$41

Math Fun

$31

$4.35

2. Reading Fun & Analysis

$52

Analysis

$46

$5.00

3. All three

$73

Using the stand-alone method with selling price as the weight for revenue allocation, what amount of revenue will be allocated to Math Fun in the package that contains all three products? (Do not round any intermediary calculations.)

A.

$29.39

B.

$23.33

C.

$16.51

D.

$31.00
Answer #1

Weight of math fun taking selling price into account

= 31/(20+31+46) = 0.31959

Amount of revenue allocated

= 73*0.31959

= 23.33

Option B is the answer